# 00footnotetext: Received xxxx Month xxxx

### v1

@@ Abstract
We present a measurement of the integrated luminosity of $e^{+}e^{-}$ collision data collected with the BESIII detector at the BEPCII collider at a center-of-mass energy of $E_{\rm cm}=3.773$ GeV. The integrated luminosities of the data sets taken from December 2021 to June 2022, from November 2022 to June 2023, and from October 2023 to February 2024 are determined to be $(4.995\pm 0.019)$ fb-1, $(8.157\pm 0.031)$ fb-1, and $(4.191\pm 0.016)$ fb-1, respectively, by analyzing large angle Bhabha scattering events. The uncertainties are dominated by systematic effects and the statistical uncertainties are negligible. Our results provide essential input for future analyses and precision measurements.
@@ 1 Introduction
where $\mathcal{L}$ denotes the integrated luminosity of the data set, $\sigma_{e^{+}e^{-}\to X}$ is the cross section for the process $e^{+}e^{-}\to X$, and $E_{\rm cm}$ is the center-of-mass energy. In principle, any process with a known cross section can be used to determine luminosity. However, Quantum Electrodynamics (QED) processes are advantageous due to their high production rates, simple final state topologies, and cross sections known with high theoretical precision.
Datasets at $E_{\rm cm}=3.773$ GeV were collected by the BESIII detector at the BEPCII collider from December 2021 to June 2022 (DATA I), from November 2022 to June 2023 (DATA II), and from October 2023 to February 2024 (DATA III) in order to systematically investigate the properties of $\psi(3770)$ and $D$ meson production and decays. Precisely determining the luminosity of this dataset is important for various purposes. It is essential for measuring the cross section of $e^{+}e^{-}\to\psi(3770)\to D\bar{D}$, calculating the normalization factors in strong-phase measurements of $D^{0}$ decays [1, 2], and reducing the systematic uncertainty of analyses using the single tag method. Additionally, the luminosity is used to normalize the Monte Carlo (MC) sample size and to estimate continuous background in the $\psi(3686)$ dataset. This paper presents a measurement of the integrated luminosity in DATA I, DATA II, and DATA III using large angle Bhabha scattering events. This dataset is currently the largest collection of $e^{+}e^{-}$ collision data at the $\psi(3770)$ resonance peak in the world.
@@ 2 BESIII detector and Monte Carlo simulation
The BESIII detector [3] records symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring [4] in the center-of-mass energy range from 2.0 to 4.95 GeV, with a peak luminosity of $1.1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $E_{\rm cm}=3.773\;\text{GeV}$ in 2023. BESIII has collected large data samples in this energy region [5, 6, 7]. The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a plastic scintillator time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identification modules interleaved with steel. The charged-particle momentum resolution at $1\leavevmode\nobreak\ {\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the TOF barrel region is 68 ps, while that in the end cap region is 60 ps [8].
Simulated data samples produced with a geant4-based [9] Monte Carlo (MC) package, which includes the geometric description of the BESIII detector and the detector response, are used to determine detection efficiencies and to estimate backgrounds. The simulation models the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc [10]. The inclusive MC sample includes the production of $D\bar{D}$ pairs (including quantum coherence for the neutral $D$ channels), the non-$D\bar{D}$ decays of the $\psi(3770)$, the ISR production of the $J/\psi$ and $\psi(3686)$ states, and the continuum processes incorporated in kkmc [10].
All particle decays are modelled with evtgen [11] using branching fractions either taken from the Particle Data Group [12], when available, or otherwise estimated with lundcharm [13]. Final state radiation (FSR) from charged final state particles is incorporated using the photos package [14].
The Bhabha signal events ($e^{+}e^{-}\to(\gamma)e^{+}e^{-}$) are simulated with the Babayaga@NLO generator [15, 16, 17, 18]. The configuration parameters for the Babayaga@NLO generator are listed in Table LABEL:optionMC. To account for beam energy fluctuations during data taking, a run-by-run calibration of $E_{\rm cm}$ has been performed and the calibrated $E_{\rm cm}$ is used as the center-of-mass energy in the generator. The duration of each run is typically one hour. The calibration of $E_{\rm cm}$, which takes advantage of the small uncertainty in the known $D$ mass $m_{D}$ [12] and the excellent resolution of the MDC and EMC, is carried out using the following equations:
@@ 3 Method
The integrated luminosity of data is usually measured with three QED processes: $e^{+}e^{-}\to(\gamma)e^{+}e^{-}$, $e^{+}e^{-}\to(\gamma)\gamma\gamma$ and $e^{+}e^{-}\to(\gamma)\mu^{+}\mu^{-}$ [^1] The symbol “($\gamma$)” represents the possible presence of photon(s) resulting from ISR or FSR. However, the uncertainties associated with reconstructing $\gamma$ and $\mu^{+(-)}$ particles are larger than those of $e^{+(-)}$. Moreover, Bhabha events at small angles, i.e., in the direction of the $e^{+}e^{-}$ beam, predominantly interact with the end cap region of the detector. This region exhibits more gaps and lower resolution compared to the barrel region. The simulation at small angles is also less precise. Given that large angle Bhabha events provide negligible statistical uncertainty, we exclusively utilize them to measure the integrated luminosity of the data set.  [^2] The large angle region is defined as $|\cos\theta|<0.83$, where $\theta$ is the polar angle of the final state electron (positron) relative to the beam direction.
@@
In this equation, the index $i$ indicates different runs and the integrated luminosity of the data set is obtained by summing over these runs. The variables $N_{i}^{\rm obs}$, $\sigma_{i}$, $\epsilon_{i}$, $\eta$, and $\epsilon^{\rm trig}$ represents the number of observed Bhabha events, the production cross section of the Bhabha process, the detection efficiency, the contamination ratio, and the trigger efficiency for collecting events in the on-line data acquisition system, respectively. The Babayaga@NLO generator [15, 16, 17, 18] is employed, using the $E_{\rm cm}$ calibrated on a run-by-run basis, to calculate the cross section, generate signal MC events for the process $e^{+}e^{-}\to(\gamma)e^{+}e^{-}$, and estimate the detection efficiency.
@@ 4.1 Event selection
Candidate Bhabha events are required to have exactly two oppositely charged tracks detected in the MDC. Each track is required to satisfy a distance of closest approach to the interaction point of less than 10 cm along the $z$ axis (the symmetry axis of the MDC) and less than 1 cm in the transverse plane. Additionally, the candidate track must lie within the polar angle region $|\rm cos\theta|<0.8$ to ensure interaction with the barrel of the EMC.
To suppress the $e^{+}e^{-}\to(\gamma)\mu^{+}\mu^{-}$ background, the deposited energy in the EMC ($E_{\rm EMC}$), as shown in Fig. 4.1, must fall within the range of $2.51.0<E_{\rm EMC}<$ 2.5 GeV. Furthermore, the sum of the momenta of the two tracks, as shown in Fig. 4.1, is required to be larger than $0.9\times E_{\rm cm}$ to suppress background events from processes involving a $J/\psi$ in the final state, e.g., $e^{+}e^{-}\to(\gamma)J/\psi$, $e^{+}e^{-}\to(\gamma)\psi(3686)\to(\gamma)J/\psi X$, and $e^{+}e^{-}\to\psi(3770)\to(\gamma)J/\psi X$ ($J/\psi\to e^{+}e^{-}$ and $X=\pi^{0}\pi^{0},\eta,\pi^{0}$ or $\gamma\gamma$). To further eliminate the background from energetic cosmic rays, the momenta of each track must be less than $E_{\rm cm}/2+0.30$ GeV.
The two oppositely charged tracks in the candidate Bhabha scattering events are expected to be back-to-back. However, due to the presence of a magnetic field, their trajectories are bent, resulting in their respective shower clusters in the $xy$-plane of the EMC not being back-to-back. To quantify this angular difference, the variable $\delta\phi$ is defined as $|\phi_{1}-\phi_{2}|-180^{\circ}$, where $\phi_{1}$ and $\phi_{2}$ are the azimuthal angles of the two clusters in the EMC. The distribution of $\delta\phi$ is illustrated in Fig. 4.1 and a requirement of $5^{\circ}<|\delta\phi|<40^{\circ}$ is imposed. This requirement effectively eliminates the background from the $e^{+}e^{-}\to(\gamma)\gamma\gamma$ process.
@@
The distribution of $E^{e^{+}}_{\rm EMC}$ versus $E^{e^{-}}_{\rm EMC}$ from a subset of the data sample. The cluster concentrated at the upper right corner is the Bhabha signal and the small cluster at the lower left corner is the $e^{+}e^{-}\to(\gamma)\mu^{+}\mu^{-}$ background.
@@
The distribution of $p_{e^{+}}+p_{e^{-}}$ from a subset of the data sample. The pink arrow indicates the $0.9\times E_{\rm cm}$ requirement.
@@
The distribution of $\delta\phi$ between the selected $e^{+}$ and $e^{-}$ tracks from a subset of the data sample. The pink arrows indicate the $5^{\circ}<|\delta\phi|<40^{\circ}$ requirement.
@@ 4.2 Background estimation
The residual background events are from various sources such as the ISR production of $J/\psi$ and $\psi(3686)$, $e^{+}e^{-}\to(\gamma)\gamma\gamma$, $e^{+}e^{-}\to\psi(3770)\to D\bar{D}$, $e^{+}e^{-}\to\psi(3770)\to$ non-$D\bar{D}$, and $e^{+}e^{-}\to$ continuum light hadron processes. These background contributions are estimated by analyzing background MC events, including $e^{+}e^{-}\to(\gamma)J/\psi$, $e^{+}e^{-}\to(\gamma)\psi(3686)$, $e^{+}e^{-}\to(\gamma)\gamma\gamma$, $e^{+}e^{-}\to(\gamma)\mu^{+}\mu^{-}$, $e^{+}e^{-}\to(\gamma)\tau^{+}\tau^{-}$, $e^{+}e^{-}\to\psi(3770)\to D^{+}D^{-}$, $e^{+}e^{-}\to\psi(3770)\to D^{0}\bar{D}^{0}$, $e^{+}e^{-}\to\psi(3770)\to$ non-$D\bar{D}$, and $e^{+}e^{-}\to q\bar{q}$. The contamination rate for the candidate Bhabha events is calculated as $\eta=N^{\rm MC}_{\rm bkg}/(N^{\rm MC}_{\rm BB}+N^{\rm MC}_{\rm bkg})$, where $N^{\rm MC}_{\rm BB}$ and $N^{\rm MC}_{\rm bkg}$ are the numbers of Bhabha and background MC events that satisfy the selection criteria, respectively. These numbers are normalized according to individual cross sections. The resulting contamination rate is $\eta=3\times 10^{-4}$.
@@ 4.4 Integrated luminosities
Inserting the number of observed Bhabha events, the trigger efficiency, the contamination rate (Sec. 4.2), the detection efficiency, and the cross sections within $|\cos\theta|<0.83$ (Sec 4.3) into Eq. (3), the integrated luminosities are determined to be $(4.995\pm 0.019)$ fb-1 for DATA I, $(8.157\pm 0.031)$ fb-1 for DATA II, and $(4.191\pm 0.016)$ fb-1 for DATA III, where the statistical uncertainties are negligible and the systematic uncertainties will be discussed in the next subsection.
@@ 4.5 Systematic uncertainty
The systematic uncertainty arising from the MDC information, which includes the MDC tracking efficiency and the momentum requirement, is determined to be 0.29% by comparing the integrated luminosities measured with and without the MDC information. The systematic uncertainty associated with the $E_{\rm EMC}$ requirements is estimated by varying the requirement from 0.8 to 0.9 or 1.1 GeV. This variation results in a change of 0.16% in the luminosity for both tracks. To estimate the systematic uncertainty caused by the $\cos\theta$ requirement, the integrated luminosity is determined with $|\cos\theta|<0.75$ or 0.70, and the difference from the standard selection of $|\cos\theta|<0.80$ is found to be 0.13% for both tracks, which is assigned as the corresponding systematic uncertainty. The uncertainty due to the $\delta\phi$ signal region selection is estimated to be 0.02% by comparing the integrated luminosities obtained by changing the lower limit from $5^{\circ}$ to $0^{\circ}$, or the higher limit from $40^{\circ}$ to $20^{\circ}$ or $30^{\circ}$. The uncertainty arising from the trigger efficiency [20] and the theoretically calculated cross section using the Babayaga@NLO generator [17] are both 0.1%. The uncertainties from the MC statistics and the contamination rate are negligible.

### v2

@@ Abstract
We present a measurement of the integrated luminosity of $e^{+}e^{-}$ collision data collected with the BESIII detector at the BEPCII collider at a center-of-mass energy of $E_{\rm cm}=3.773$ GeV. The integrated luminosities of the data sets taken from December 2021 to June 2022, from November 2022 to June 2023, and from October 2023 to February 2024 are determined to be $4.995\pm 0.019$ fb-1, $8.157\pm 0.031$ fb-1, and $4.191\pm 0.016$ fb-1, respectively, by analyzing large angle Bhabha scattering events. The uncertainties are dominated by systematic effects and the statistical uncertainties are negligible. Our results provide essential input for future analyses and precision measurements.
@@ 1 Introduction
where $\mathcal{L}$ denotes the integrated luminosity of the data set, $\sigma_{e^{+}e^{-}\to X}$ is the cross section for the process $e^{+}e^{-}\to X$, and $E_{\rm cm}$ is the center-of-mass energy. In principle, any process with a known cross section can be used to determine luminosity. However, Quantum Electrodynamics (QED) processes are advantageous due to their high production rates, simple final state topologies, and cross sections that are known with high theoretical precision.
Datasets at $E_{\rm cm}=3.773$ GeV were collected by the BESIII detector at the BEPCII collider from December 2021 to June 2022 (DATA I), from November 2022 to June 2023 (DATA II), and from October 2023 to February 2024 (DATA III) in order to systematically investigate the properties of $\psi(3770)$ and $D$ meson production and decays. Precisely determining the luminosity of this dataset is important for various purposes. It is essential for measuring the cross section of $e^{+}e^{-}\to\psi(3770)\to D\bar{D}$, calculating the normalization factors in strong-phase measurements of $D^{0}$ decays [1, 2], and reducing the systematic uncertainty of analyses using the single tag method [3]. Additionally, the luminosity is used to normalize the Monte Carlo (MC) sample size and to estimate continuous background in the $\psi(3686)$ dataset [4, 5]. This paper presents a measurement of the integrated luminosity in DATA I, DATA II, and DATA III using large angle Bhabha scattering events. This dataset is currently the world’s largest collection of $e^{+}e^{-}$ collision data at the $\psi(3770)$ resonance peak.
@@ 2 BESIII detector and Monte Carlo simulation
The BESIII detector [6] records energy-symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring [7] in the $E_{\rm cm}$ range from 2.0 to 4.95 GeV, with a peak luminosity of $1.1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $E_{\rm cm}=3.773\;\text{GeV}$ in 2023. BESIII has collected large data samples in this energy region [8, 9, 10]. The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a plastic scintillator time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identification modules interleaved with steel. The charged-particle momentum resolution at $1\leavevmode\nobreak\ {\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the TOF barrel region is 68 ps, while that in the end cap region is 60 ps [11].
Simulated data samples produced with a geant4-based [12] Monte Carlo (MC) package, which includes the geometric description of the BESIII detector and the detector response, are used to determine detection efficiencies and to estimate backgrounds. The QED processes are simulated with the Babayaga@NLO generator [13, 14, 15, 16]. The width of the $\psi(3770)$, beam energy spread, initial state radiation (ISR), and final state radiation (FSR) are considered in the simulation. The configuration parameters of the Bhabha signal MC ($e^{+}e^{-}\to(\gamma)e^{+}e^{-}$) are listed in Table LABEL:optionMC. As for other processes, the simulation models the beam energy spread and ISR in the $e^{+}e^{-}$ annihilations with the generator kkmc [17]. The inclusive MC sample is used to simulate all possible processes of $e^{+}e^{-}$ collision, including the production of $D\bar{D}$ pairs (including quantum coherence for the neutral $D$ channels), the non-$D\bar{D}$ decays of the $\psi(3770)$, the ISR production of the $J/\psi$ and $\psi(3686)$ states, and the continuum processes incorporated in kkmc [17].
All particle decays are modelled with evtgen [18] using branching fractions either taken from the Particle Data Group [19], when available, or otherwise estimated with lundcharm [20]. FSR from charged final state particles is incorporated using the photos package [21].
In the simulation above, the run-by-run based calibration is applied, taking into consideration the beam energy fluctuations and other time-dependent effects. The duration of each run is typically one hour. The calibration of $E_{\rm cm}$, which takes advantage of the small uncertainty in the known $D$ mass $m_{D}$ [19] and the excellent resolution of the MDC and EMC, is carried out using the following equations:
@@ 3 Method
The integrated luminosity of data is usually measured with three QED processes: $e^{+}e^{-}\to(\gamma)e^{+}e^{-}$, $e^{+}e^{-}\to(\gamma)\gamma\gamma$ and $e^{+}e^{-}\to(\gamma)\mu^{+}\mu^{-}$. The symbol “($\gamma$)” represents the possible presence of photon(s) resulting from ISR or FSR. Since the uncertainties associated with reconstructing electrons are smaller than those of photons and muons, and the statistical uncertainty of Bhabha events are almost negligible, only Bhabha event is used to measure luminosity. Furthormore, Bhabha events at small angles, i.e., in the direction of the $e^{+}e^{-}$ beam, predominantly interact with the end cap region of the detector. This region exhibits more gaps and worse resolution compared to the barrel region. The generator simulation at small angles is also less precise. Given that large angle Bhabha events provide a negligible statistical uncertainty, we exclusively utilize them to measure the integrated luminosity of the data set. The large angle region is defined as $|\cos\theta|<0.83$, where $\theta$ is the polar angle of the final state electron (positron) relative to the beam direction.
@@
In this equation, the index $i$ indicates different runs, and the integrated luminosity of the data set is obtained by summing over these runs. The variables $N_{i}^{\rm obs}$, $\sigma_{i}$, $\epsilon_{i}$, $\eta$, and $\epsilon^{\rm trig}$ represent the number of observed Bhabha-event candidates, the production cross section of the Bhabha process, the detection efficiency, the contamination rate, and the trigger efficiency for collecting events in the on-line data acquisition system, respectively. The Babayaga@NLO generator [13, 14, 15, 16] is employed, using the $E_{\rm cm}$ calibrated on a run-by-run basis, to calculate the cross section, generate signal MC events for the process $e^{+}e^{-}\to(\gamma)e^{+}e^{-}$, and estimate the detection efficiency.
@@ 4.1 Event selection
Candidate Bhabha events are required to have exactly two oppositely charged tracks detected in the MDC. Each track is required to satisfy a distance of closest approach to the interaction point of less than 10 cm along the $z$ axis (the symmetry axis of the MDC) and less than 1 cm in the transverse plane. Additionally, the each candidate track must lie within the polar angle region $|\rm cos\theta|<0.8$ to ensure interaction with the barrel of the EMC.
To suppress the $e^{+}e^{-}\to(\gamma)\mu^{+}\mu^{-}$ background, the deposited energy $E_{\rm EMC}$ in the EMC by each track, shown in Fig. 4.1, must fall within the range $2.51.0<E_{\rm EMC}<$ 2.5 GeV. Furthermore, the sum of the momenta of the two tracks, shown in Fig. 4.1, is required to be larger than $0.9\times E_{\rm cm}$ to suppress background events, which mainly from processes involving a $J/\psi$ in the final state, e.g., $e^{+}e^{-}\to(\gamma)J/\psi$, $e^{+}e^{-}\to(\gamma)\psi(3686)\to(\gamma)J/\psi X$, and $e^{+}e^{-}\to\psi(3770)\to(\gamma)J/\psi X$. To eliminate the background from energetic cosmic rays, the momentum of each track must be less than $E_{\rm cm}/2+0.30$ GeV.
The two oppositely charged tracks in the candidate Bhabha scattering events are produced to be back-to-back. However, due to the presence of a magnetic field, their trajectories are bent, resulting in their respective shower clusters in the $xy$-plane of the EMC not being back-to-back. To quantify this angular difference, the variable $\delta\phi$ is defined as $|\phi_{1}-\phi_{2}|-180^{\circ}$, where $\phi_{1}$ and $\phi_{2}$ are the azimuthal angles of the two clusters in the EMC. The distribution of $\delta\phi$ is illustrated in Fig. 4.1 and a requirement of $5^{\circ}<|\delta\phi|<40^{\circ}$ is imposed. This requirement effectively eliminates the background from the $e^{+}e^{-}\to(\gamma)\gamma\gamma$ process.
@@
The distribution of $E^{e^{+}}_{\rm EMC}$ versus $E^{e^{-}}_{\rm EMC}$ from a subset of the data sample. The cluster concentrated at the upper right corner is the Bhabha signal and the small cluster at the lower left corner is the $e^{+}e^{-}\to(\gamma)\mu^{+}\mu^{-}$ background. The horizontal and vertical bands are caused by ISR and FSR in Bhabha and $e^{+}e^{-}\to(\gamma)\mu^{+}\mu^{-}$ events.
@@
The distribution of $p_{e^{+}}+p_{e^{-}}$ from a subset of the data sample. The background is not visible due to the very low contamination rate of $\eta=3\times 10^{-4}$. The pink arrow indicates the $>0.9\times E_{\rm cm}$ requirement.
@@
The distribution of $\delta\phi$ between the selected $e^{+}$ and $e^{-}$ tracks from a subset of the data sample. The background is not visible due to the very low contamination rate of $\eta=3\times 10^{-4}$. The pink arrows indicate the $5^{\circ}<|\delta\phi|<40^{\circ}$ requirement.
@@ 4.2 Background estimation
The residual background events are from various sources, such as the ISR production of $J/\psi$ and $\psi(3686)$, $e^{+}e^{-}\to(\gamma)\gamma\gamma$, $e^{+}e^{-}\to\psi(3770)\to D\bar{D}$, $e^{+}e^{-}\to\psi(3770)\to$ non-$D\bar{D}$, and $e^{+}e^{-}\to$ continuum processes. These background contributions are estimated by analyzing the corresponding background MC events. The contamination rate for the candidate Bhabha events is calculated as $\eta=N^{\rm MC}_{\rm bkg}/(N^{\rm MC}_{\rm BB}+N^{\rm MC}_{\rm bkg})$, where $N^{\rm MC}_{\rm BB}$ and $N^{\rm MC}_{\rm bkg}$ are the numbers of Bhabha and background MC events that satisfy the selection criteria, respectively. These numbers are normalized according to individual cross sections. The resulting contamination rate is $\eta=3\times 10^{-4}$.
@@ 4.4 Integrated luminosities
Inserting the number of observed Bhabha events, the trigger efficiency, the contamination rate (Sec. 4.2), the detection efficiency, and the cross sections within $|\cos\theta|<0.83$ (Sec 4.3) into Eq. (3), the integrated luminosities are determined to be $4.995\pm 0.019$ fb-1 for DATA I, $8.157\pm 0.031$ fb-1 for DATA II, and $4.191\pm 0.016$ fb-1 for DATA III, where the statistical uncertainties are negligible and the systematic uncertainties will be discussed in the next subsection.
@@ 4.5 Systematic uncertainty
The systematic uncertainty arising from the MDC information, which includes the MDC tracking efficiency and the momentum requirement, is determined to be 0.29% by comparing the integrated luminosities measured with and without the MDC information. The distribution of $\delta\phi$ without the MDC information applied is shown in Fig. 4.5, and the efficiency increases by about 5%. The dominant background is $e^{+}e^{-}\to(\gamma)\gamma\gamma$, which is at the 0.15% level. The systematic uncertainty associated with the $E_{\rm EMC}$ requirements is estimated by varying the requirement from 1.0 to 0.9 or 1.1 GeV. This variation results in a change of 0.16% in the luminosity for both tracks. To estimate the systematic uncertainty caused by the $\cos\theta$ requirement, the integrated luminosity is determined with $|\cos\theta|<0.75$ or 0.70, and the difference from the standard selection of $|\cos\theta|<0.80$ is found to be 0.13% for both tracks, which is assigned as the corresponding systematic uncertainty. The uncertainty due to the $\delta\phi$ signal region selection is estimated to be 0.02% by comparing the integrated luminosities obtained by changing the lower limit from $5^{\circ}$ to $0^{\circ}$, or the higher limit from $40^{\circ}$ to $20^{\circ}$ or $30^{\circ}$. The uncertainty arising from the trigger efficiency [23] and the theoretically calculated cross section using the Babayaga@NLO generator [15, 24, 25] are both 0.1%. The uncertainties from the MC statistics and the contamination rate are negligible.
[…]
The distribution of $\delta\phi$ between the selected $e^{+}$ and $e^{-}$ tracks without MDC information from a subset of the data sample.